OpenStudy (anonymous):

Solve the following equations x^a^x=a^x^a, knowing that a<1 and a>0. I love this exercise!

OpenStudy (anonymous):

x^(ax)= a^(ax) ==> ln [x^(ax)]= ln [a^(ax)] ===> ax *ln x= ax ln a ===> ln x = ln a ===> x=a

OpenStudy (anonymous):

Nope. BEWARE! It is not x^(ax). It is x^a^x. NOT THE SAME THING!!I wouldn't post a easy question! \[x^{a^{x}}=a^{x^{a}}\]

OpenStudy (anonymous):

it's really a matter of notations writing method online, you maight be right. Thank you.

OpenStudy (anonymous):

in your case: x^(a^x) = a^(x^a) , this is the conventional way to write the experession withouth math editor

OpenStudy (anonymous):

not as been posted at the begning

OpenStudy (anonymous):

My mistake :)

OpenStudy (anonymous):

x=a is still a solution. it should be obvious that: \[a^{a^a}=a^{a^a}\]

OpenStudy (anonymous):

Yes, it is true! x=a is a solution, indeed, but one has to show that it is unique. And this is the nice part of the problem....

OpenStudy (anonymous):

ok, great then to make such a momentum.

OpenStudy (anonymous):

Good luck with it!